OpenStudy (anonymous):

Samantha is making a beaded necklaces using 54 glass beads and 18 clay beads. She wants each necklace to have the same number of beads, but each necklace will have only one type of bead. If she puts the greatest possible number of beads on each necklace, how many necklace can she make?

OpenStudy (anonymous):

This question is worded soo unclearly. Im trying to wrap my brain around it

OpenStudy (anonymous):

maxm beads can be 18 number of necklaces with glass beads= 54/18=3 number of necklaces with clay beads=18/18=1

OpenStudy (anonymous):

I just copied the question for a California textbook its my last homework question

OpenStudy (anonymous):

does it give you choices?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its 4

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome